# RSCC: A Large-Scale Remote Sensing Change Caption Dataset for Disaster Events

### v1

@@ 1 Introduction
Existing remote sensing image-text datasets often focus on single-snapshot imagery and lack the temporal details vital for understanding dynamic events, particularly in disaster-related scenarios, As shown in Table 1. Although there are multi-temporal datasets (e.g ., fMoW (Christie et al., 2018), SpaceNet 7 (Van Etten et al., 2021), S2Looking (Shen et al., 2021), QFabric (Verma et al., 2021) and SpaceNet 8 (Hänsch et al., 2022)), none of them provide rich textual descriptions of how scenes change over time. However, their potential for disaster-specific temporal analysis remains untapped due to the absence of high-quality bi-temporal datasets with detailed textual annotations. Existing remote sensing datasets either focus on generic land-use changes or provide short captions lacking disaster context. For instance, LEVIR-CC (Liu et al., 2022b) annotates urban development but omits disaster-specific details, while Dubai-CCD (Hoxha et al., 2022b) offers brief descriptions without capturing nuanced damage levels or infrastructure transformations.
@@ 4.1 Experiment Settings
For remote sensing change captioning, we benchmark moderate size open-sourced MLLMs (less than 13B parameters) that supports multi-image inputs, including LLaVA-NeXT-Interleave (Li et al., 2024b), xGen-MM [^2] https://hugggingfae.co/Salesforce/xgen-mm-phi3-mini-instruct-interleave-r-v1.5 (BLIP-3) (Xue et al., 2024), LLaVA-OneVision (Li et al., 2024a), Qwen2-VL (Wang et al., 2024c),Pixtral (Mistral AI, 2024) Phi-4-Multimodal (Microsoft, 2025) Kimi-VL (Kimi Team, 2025) and InternVL 3 (Zhu et al., 2025). We also add two specialized remote sensing change captioning models namely TEOChat (Irvin et al., 2024) and CCExpert (Wang et al., 2024e).
@@ 4.3 Human Preference Study
While the language metrics can be biased, we ask experts to vote the best caption from two anonymous model output given the bi-temporal image pairs along with human labeled building damage masks from xBD dataset Gupta et al. (2019). Results (Figure 4) reveal QvQ-Max (ground truth change captions) consistently outperformed all baselines, achieving win rates ranging from 80.7% (against InternVL3) to 99.0% (against CCExpert). While strong baselines like InternVL3 (19.3% wins) and mid-tier models (e.g., Pixtral [14.6%], Kimi-VL [12.8%]) showed moderate performance, our captions demonstrated superior accuracy in capturing fine-grained environmental changes critical for disaster response. Weak-performing multimodal baselines (LLaVA-Interleave [5.2%], Phi-4-MM [4.9%]) highlighted limitations in handling complex spatiotemporal reasoning, suggesting QvQ-Max’s quantization-aware training and dynamic context adaptation mechanisms enhance generalization. These findings validate QvQ-Max as a state-of-the-art solution for vision-language tasks in remote sensing.
@@ 5.2 Remote Sensing Change Caption Dataset
Datasets that combine temporal and vision-language elements play a crucial role in training models to comprehend and merge temporal dynamics with linguistic information Liu et al. (2024a). Unlike the VQA dataset Irvin et al. (2024); Elgendy et al. (2024) that can be easily formatted through mask labels, the common practice of remote sensing change caption dataset is to further annotate existing change detection dataset with 5 sentences each image pairs, such as Dubai-CCD (Hoxha et al., 2022b), LEVIR-CC (Liu et al., 2023) and WHU-CDC (Shi et al., 2024). Given the generalization capability of commercial MLLMs, Wang et al.  Wang et al. (2024e) leveraged GPT-4o OpenAI (2024), using the explicit information provided by the change masks to generate detailed change descriptions.
@@ A.2 More Results
In Scenario (c), ground truth provides the most accurate, complete, and factually consistent description. It captures the catastrophic scale of destruction ("every structure reduced to rubble," "barren landscape"), explicitly mentions debris and empty streets, and aligns with typical patterns of severe wind-driven disasters (e.g., hurricanes or tornadoes). While it does not specify the disaster type, its focus on observable damage patterns (total structural collapse, vegetation loss) adheres strictly to visual evidence. Other captions either misinterpret the scene (Phi-4-MM, CCExpert), lack detail (TEOChat), or omit critical damage indicators (Qwen2-VL).
@@ A.4 Author Statement and Data License
Data License Confirmation: The dataset is released under the [CC-BY-4.0], which permits unrestricted use, distribution, and reproduction in any medium, provided the original work is properly cited.

### v2

@@ 1 Introduction
Existing remote sensing image-text datasets often focus on single-snapshot imagery and lack the temporal details vital for understanding dynamic events, particularly in disaster-related scenarios, As shown in Table 1. Although there are multi-temporal datasets (e.g ., fMoW (Christie et al., 2018), SpaceNet 7 (Van Etten et al., 2021), S2Looking (Shen et al., 2021), QFabric (Verma et al., 2021) and SpaceNet 8 (Hänsch et al., 2022)), none of them provide rich textual descriptions of how scenes change over time. However, their potential for disaster-specific temporal analysis remains untapped due to the absence of high-quality bi-temporal datasets with detailed textual annotations. Existing remote sensing datasets either focus on generic land-use changes or provide short captions lacking disaster context. For instance, LEVIR-$C$ (Liu et al., 2022b) annotates urban development but omits disaster-specific details, while Dubai-CCD (Hoxha et al., 2022b) offers brief descriptions without capturing nuanced damage levels or infrastructure transformations.
@@ 4.1 Experiment Settings
For remote sensing change captioning, we benchmark moderate size open-sourced MLLMs (less than 13B parameters) that supports multi-image inputs, including LLaVA-NeXT-Interleave (Li et al., 2024b), xGen-MM [^2] https://hugggingfae.co/Salesforce/xgen-$m$-phi3-mini-instruct-interleave-r-v1.5 (BLIP-3) (Xue et al., 2024), LLaVA-OneVision (Li et al., 2024a), Qwen2-VL (Wang et al., 2024c),Pixtral (Mistral AI, 2024) Phi-4-Multimodal (Microsoft, 2025) Kimi-VL (Kimi Team, 2025) and InternVL 3 (Zhu et al., 2025). We also add two specialized remote sensing change captioning models namely TEOChat (Irvin et al., 2024) and CCExpert (Wang et al., 2024e).
@@ 4.3 Human Preference Study
While the language metrics can be biased, we ask experts to vote the best caption from two anonymous model output given the bi-temporal image pairs along with human labeled building damage masks from xBD dataset Gupta et al. (2019). Results (Figure 4) reveal QvQ-Max (ground truth change captions) consistently outperformed all baselines, achieving win rates ranging from 80.7% (against InternVL3) to 99.0% (against CCExpert). While strong baselines like InternVL3 (19.3% wins) and mid-tier models (e.g., Pixtral [14.6%], Kimi-VL [12.8%]) showed moderate performance, our captions demonstrated superior accuracy in capturing fine-grained environmental changes critical for disaster response. Weak-performing multimodal baselines (LLaVA-Interleave [5.2%], Phi-4-$M$ [4.9%]) highlighted limitations in handling complex spatiotemporal reasoning, suggesting QvQ-Max’s quantization-aware training and dynamic context adaptation mechanisms enhance generalization. These findings validate QvQ-Max as a state-of-the-art solution for vision-language tasks in remote sensing.
@@ 5.2 Remote Sensing Change Caption Dataset
Datasets that combine temporal and vision-language elements play a crucial role in training models to comprehend and merge temporal dynamics with linguistic information Liu et al. (2024a). Unlike the VQA dataset Irvin et al. (2024); Elgendy et al. (2024) that can be easily formatted through mask labels, the common practice of remote sensing change caption dataset is to further annotate existing change detection dataset with 5 sentences each image pairs, such as Dubai-CCD (Hoxha et al., 2022b), LEVIR-$C$ (Liu et al., 2023) and WHU-CDC (Shi et al., 2024). Given the generalization capability of commercial MLLMs, Wang et al.  Wang et al. (2024e) leveraged GPT-4o OpenAI (2024), using the explicit information provided by the change masks to generate detailed change descriptions.
@@ A.2 More Results
In Scenario (c), ground truth provides the most accurate, complete, and factually consistent description. It captures the catastrophic scale of destruction ("every structure reduced to rubble," "barren landscape"), explicitly mentions debris and empty streets, and aligns with typical patterns of severe wind-driven disasters (e.g., hurricanes or tornadoes). While it does not specify the disaster type, its focus on observable damage patterns (total structural collapse, vegetation loss) adheres strictly to visual evidence. Other captions either misinterpret the scene (Phi-4-$M$, CCExpert), lack detail (TEOChat), or omit critical damage indicators (Qwen2-VL).
@@ A.4 Author Statement and Data License
Data License Confirmation: The dataset is released under the [$C$-BY-4.0], which permits unrestricted use, distribution, and reproduction in any medium, provided the original work is properly cited.